How do i work out a problem where you have two consecutive even integers, such that five times the first integer is equal to four times the second integer?
Answer by ewatrrr(24785)


Hi
two consecutive even integers
Let x and (x+2)represent the smaller and larger respectively
Question states***
5x = 4(x+2)
Solving for x
x = 8 The two consecutive even integers are 8,10
CHECKING our Answer***
5*8 = 4*10 = 40